Fermi National Laboratory
 

Accelerator Update for April 10, 2009 - April 13, 2009

 

The Integrated Luminosity 4/6/09 to 4/13/09 was 27.08 pb -1.

Friday April13

The day shift began with Operations monitoring the Tevatron (TeV) store 6965 (7.7E31) and the Recycler stash (260.9E10), with the Antiproton Source stacking (26.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

The store was terminated at 9:45 AM and HEP shot setup followed.

Store 6966 was colliding by 11:15 AM with an average initial luminosity of 29.0E31.

MTest ceased data taking for the day at 4:45 PM.

During the evening shift, the I- Source high voltage power supply tripped and an investigation ensued.

At 6:00 PM, A:IB, the Accumulator bend bus power supply, ground faulted. The supply was turned on and stacking resumed at 6:20 PM.

At 6:25 PM, the store 6966 was aborted due to the TeV sector A3 VCB opening. EE Support personnel came in and began replacement of the VCB.

Linac personnel replaced the anode power supply and Proton Source beam was available by 6:30 PM.

A:IB tripped off again at 10:00 PM and Pbar personnel came in and investigated.

Saturday April 11

The midnight shift began with the TeV in recovery, with Operations monitoring the stash (203.9E10), with Pbar off for A:IB repairs, and with NuMI and MiniBooNE taking beam.

A:IB ground fault sensor was swapped with D:IB and stacking resumed by 1:30 AM.

TeV repairs were complete by 2:05 AM.

A:IB faulted off again at 2:15 AM. Pbar personnel accessed their enclosures for a ground fault hunt, which was not conclusive.

HEP shot setup began at 4:05 AM. Store 6969 was colliding by 5:30 AM with an average initial luminosity of 21.8E31.

Spill was established to MTest at 5:40 AM.

A:IB ground faults were numerous on the day shift. Ops transferred antiprotons to the Accumulator when the ground current would fluctuate.

MTest spill was disabled at 6:00 PM.

Booster RF stations 8 and 19 were turned off due to non-optimal gap envelopes, and station 1 was turned on.

A MI Physicist conducted a crystal collimator study.

Sunday April 12

The midnight shift began with Operations monitoring store 6969 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The store was terminated at 1:50 AM and shot setup followed.

Store 6970 was colliding by 3:15 AM with an average initial luminosity of 19.6E31.

Pbar personnel verified A:IB had a faulty power cable. Store 6969 aborted at 2:45 PM due to T:SF, a TeV focusing sextupole power supply, faulting off. EE Support personnel came in and investigated. The cause was determined to be a faulty mode selector switch on the regulator chassis. The switch was modified.

HEP shot setup began at 11:25 PM.

Monday April 13

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Store 6974 was colliding by 12:45 AM with an average initial luminosity of 32.3E31.

At 2:25 AM, Operations was informed of a fire alarm in the MINOS Soudan Mine. The Far Detector power was down. NuMI was set to run in horn-off mode for a special run with the Near Detector.

MTest began taking spill at 5:20 AM.

Operations reported that they monitored A:IB ground current and transferred antiprotons to Recycler prior to the power supply tripping off.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/13/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory